OpenStudy (marcelie):

Help Please !!!! How do i solve this ????? For the following exercises, find the point of intersection of each pair of lines if it exists. If it does not exist, indicate that there is no point of intersection. y =3/4 x + 1 −3x + 4y = 12

OpenStudy (anonymous):

slopes are same ,so lines are parallel.

OpenStudy (marcelie):

i set the equations equal to each ? y= 3/4x+1 y=3/4x+3 3/4x+1= 3/4x+3 -3/4x -3/4x 1= 3 ? so does that mean theres no intersection ?

OpenStudy (marcelie):

I know they are parallel so the answer doesnt make it an intersection ?

OpenStudy (anonymous):

they are parallel because slopes are same if 1=1,or 3=3 or c=c then same line.

OpenStudy (marcelie):

oh okay . so theres no intersection since they are parallel and the answer i got didnt match ?